# Rethinking Schema Linking: A Context-Aware Bidirectional Retrieval Approach for Text-to-SQL

### v1

@@ Abstract
Md Mahadi Hasan Nahid  Davood Rafiei  Weiwei Zhang  Yong Zhang University of Alberta, Canada Huawei Technologies Canada Co., Ltd., Burnaby, Canada {mnahid, drafiei}@ualberta.ca{weiwei.zhang2, yong.zhang3}@huawei.com
@@ 2 Related Work
To mitigate these issues, several recent works propose schema pruning techniques. SQL-to-Schema Yang et al. (2024) and RSL-SQL Cao et al. (2024) generates SQL over the full schema, then backtraces required tables and columns to construct a reduced schema context. Similarly, E-SQL Caferoğlu and Özgür Ulusoy (2025) enhances the original question with schema elements, helping align user intent with the database structure. Kothyari et al. (2023) introduces a retrieval-based method using hallucinated schema elements and candidate reranking, tightly coupled with SQL generation. The recent DTS-SQL Pourreza and Rafiei (2024b) framework shows that separating schema linking and SQL generation through a two-stage fine-tuning process enhances the performance of smaller models, underscoring the critical role of schema linking. Gorti et al. (2025) propose MSc-SQL, which critiques multiple sampled SQL outputs using execution feedback and metadata to select the best candidate, enabling small open-source models to rival larger closed ones.
@@ 4.1 Datasets and Baselines
We compare our method against a set of strong baselines that reflect the current state of the art in schema-aware and LLM-based Text-to-SQL systems. These include decomposed in-context learning (DIN-SQL Pourreza and Rafiei (2024a), C3SQL Dong et al. (2023)), robust schema linking (RSL-SQL Cao et al. (2024), TA-SQL Qu et al. (2024), Solid-SQL Liu et al. (2025)), direct linking with question enrichment (E-SQL Caferoğlu and Özgür Ulusoy (2025)), and multi-agent or multi-path models such as CHESS Talaei et al. (2024), MAC-SQL  Wang et al. (2024) and MCS-SQL Lee et al. (2025). Together, these baselines represent diverse and competitive strategies for addressing the challenges of schema linking and SQL generation.
@@ 4.3 Implementation Details
We use four LLMs for evaluation: GPT-4o-mini(G) OpenAI et al. (2024), DeepSeek-Chat(D) DeepSeek-AI et al. (2025), Qwen-2.5-7B-Instruct(Q) Hui et al. (2024), and Gemini-2.0-Flash(M) Team et al. (2024). Prompting strategies are adapted from prior work Talaei et al. (2024); Maamari et al. (2024). Schema linking and SQL generation are performed in two stages using structured instructions and augmented query context. Further implementation details and prompt templates are available in Appendix D and H  [^1] Code will be released upon paper acceptance.
@@ 5.2 Impact on Text-to-SQL Performance
Table 3 presents a comparison of zero-shot SQL execution accuracy on the BIRD development set between our approach and several baseline methods. Our approach achieves performance on par with strong baselines such as E-SQL and Solid-SQL. However, it is important to highlight a critical difference: unlike these methods, our framework does not incorporate any SQL query fixing, refinement, or self-consistency techniques during SQL generation. In contrast, most existing methods depend on post-generation correction pipelines to boost execution accuracy Talaei et al. (2024); Pourreza and Rafiei (2024a); Gao et al. (2024); Lee et al. (2025); Caferoğlu and Özgür Ulusoy (2025); Wang et al. (2024).
@@ 5.3 Ablation Study
5.4 Schema Reduction
We evaluate the effectiveness of our proposed methods in reducing the database schema size by measuring the percentage of tables and columns retained after applying schema linking. Table 5.3 reports the reduction rates for different variants of our method, including the full bidirectional approach, table-first, column-first, and the ablated version without augmentation. The results demonstrate that our bidirectional schema linking strategy is effective at significantly narrowing down the schema context by selecting only the most relevant tables and columns. This targeted reduction not only simplifies the reasoning space for the language model but also improves execution efficiency, especially for databases with large schemas. These findings highlight the practical advantage of incorporating schema pruning into the Text-to-SQL pipeline.
@@ Appendix A Additional Results
Plug-in with other Models.
Model Acc (%) W. Our Schema (%)
DIN-SQL 50.6 55.4
MCS-SQL 52.8 54.7
@@ Appendix B Error Analysis
These error types reflect both lexical and semantic challenges in schema linking, and they motivate future work on enhanced disambiguation, context-aware retrieval, and reasoning-aware schema grounding. Table 13 summarizes these error categories and their relative frequencies and some examples are discussed in Appendix I.
@@
Method # LLM Calls Tokens / Sample Latency (s)
CHESS Talaei et al. (2024) 78 339k 46.8
DIN-SQL Pourreza and Rafiei (2024a) 1 8.7k 1.2
SCSL Maamari et al. (2024) 78 339k 46.8
HyTCSL Maamari et al. (2024) 78 340k 46.8
TCSL Maamari et al. (2024) 2 8.7k 1.6
RSL-SQL Cao et al. (2024) 2 8.7k 1.6
TA-SQLQu et al. (2024) 1 6.1k 0.75
Ours 6 10.5k 2.6
@@
Appendix C Token-LLM-calls-latency Analysis
Appendix D Implementation Details
@@
Appendix E Design Rationale and Justification
@@
Appendix F Extended Related Work
@@
Approaches like E-SQL Caferoğlu and Özgür Ulusoy (2025) enrich questions with schema context to enable direct alignment, while RSL-SQL Cao et al. (2024) utilizes partial query execution to validate links and improve robustness.
@@
The advent of LLMs shifted the paradigm toward in-context and retrieval-based prompting. MCS-SQL Lee et al. (2025) shows that using multiple prompts and multiple-choice ranking significantly improves accuracy by broadening the SQL generation search space. Solid-SQL Liu et al. (2024a) combines schema-aware prompting with retrieval strategies to improve robustness.
@@
Appendix G Schema Linking Examples
@@
Appendix H Prompt Templates
@@
Appendix I Representative Error Case

### v2

@@ Abstract
Md Mahadi Hasan Nahid thanks: This work was conducted during an internship at Huawei Technologies Canada Co., Ltd., Burnaby, Canada.  Davood Rafiei  Weiwei Zhang  Yong Zhang University of Alberta, Canada Huawei Technologies Canada Co., Ltd., Canada {mnahid, drafiei}@ualberta.ca{weiwei.zhang2, yong.zhang3}@huawei.com
@@ 2 Related Work
To mitigate these issues, several recent works propose schema pruning techniques. SQL-to-Schema Yang et al. (2024) and RSL-SQL Cao et al. (2024) generates SQL over the full schema, then backtraces required tables and columns to construct a reduced schema context. Similarly, E-SQL caferoğlu2025esqldirectschemalinking enhances the original question with schema elements, helping align user intent with the database structure. Kothyari et al. (2023) introduces a retrieval-based method using hallucinated schema elements and candidate reranking, tightly coupled with SQL generation. The recent DTS-SQL Pourreza and Rafiei (2024b) framework shows that separating schema linking and SQL generation through a two-stage fine-tuning process enhances the performance of smaller models, underscoring the critical role of schema linking. Gorti et al. (2025) propose MSc-SQL, which critiques multiple sampled SQL outputs using execution feedback and metadata to select the best candidate, enabling small open-source models to rival larger closed ones.
@@ 4.1 Datasets and Baselines
We compare our method against a set of strong baselines that reflect the current state of the art in schema-aware and LLM-based Text-to-SQL systems. These include decomposed in-context learning (DIN-SQL Pourreza and Rafiei (2024a), C3SQL Dong et al. (2023)), robust schema linking (RSL-SQL Cao et al. (2024), TA-SQL Qu et al. (2024), Solid-SQL Liu et al. (2025)), direct linking with question enrichment (E-SQL caferoğlu2025esqldirectschemalinking), and multi-agent or multi-path models such as CHESS Talaei et al. (2024), MAC-SQL  Wang et al. (2024) and MCS-SQL Lee et al. (2025). Together, these baselines represent diverse and competitive strategies for addressing the challenges of schema linking and SQL generation.
@@ 4.3 Implementation Details
We use four LLMs for evaluation: GPT-4o-mini(G) OpenAI et al. (2024), DeepSeek-Chat(D) DeepSeek-AI et al. (2025), Qwen-2.5-7B-Instruct(Q) Hui et al. (2024), and Gemini-2.0-Flash(M) Team et al. (2024). Prompting strategies are adapted from prior work Talaei et al. (2024); Maamari et al. (2024). Schema linking and SQL generation are performed in two stages using structured instructions and augmented query context. Further implementation details and prompt templates are available in Appendix C and G. The implementation of our framework is publicly available at https://github.com/mahadi-nahid/BiSchemaLink.
@@ 5.2 Impact on Text-to-SQL Performance
Table 3 presents a comparison of zero-shot SQL execution accuracy on the BIRD development set between our approach and several baseline methods. Our approach achieves performance on par with strong baselines such as E-SQL and Solid-SQL. However, it is important to highlight a critical difference: unlike these methods, our framework does not incorporate any SQL query fixing, refinement, or self-consistency techniques during SQL generation. In contrast, most existing methods depend on post-generation correction pipelines to boost execution accuracy Talaei et al. (2024); Pourreza and Rafiei (2024a); Gao et al. (2024); Lee et al. (2025); caferoğlu2025esqldirectschemalinking; Wang et al. (2024).
@@
Plug-in with other Models.
Model Acc (%) W/ Our Schema (%)
DIN-SQL 50.6 55.4
MCS-SQL 52.8 54.7
@@ 5.3 Ablation Study
5.4 Schema Reduction and Error Analysis
We evaluate the effectiveness of our proposed methods in reducing the database schema size by measuring the percentage of tables and columns retained after applying schema linking. Table 5.3 reports the reduction rates for different variants of our method, including the full bidirectional approach, table-first, and column-first approach. The results demonstrate that our bidirectional schema linking strategy is effective at significantly narrowing down the schema context by selecting only the most relevant tables and columns. This targeted reduction not only simplifies the reasoning space for the language model but also improves execution efficiency, especially for databases with large schemas. These findings highlight the practical advantage of incorporating schema pruning into the Text-to-SQL pipeline.
Illustrative examples are presented in Appendix F, detailed error cases are discussed in Appendix H, and a comprehensive error analysis is provided in Appendix B. The appendices systematically categorize common failure modes and include representative examples from our experiments.
5.5 Token-LLM-calls-latency Analysis
Method # LLM Calls Tokens / Sample Latency (s)
CHESS Talaei et al. (2024) 78 339k 46.8
DIN-SQL Pourreza and Rafiei (2024a) 1 8.7k 1.2
SCSL Maamari et al. (2024) 78 339k 46.8
HyTCSL Maamari et al. (2024) 78 340k 46.8
TCSL Maamari et al. (2024) 2 8.7k 1.6
RSL-SQL Cao et al. (2024) 2 8.7k 1.6
TA-SQLQu et al. (2024) 1 6.1k 0.75
Ours 6 10.5k 2.6
@@ 6 Conclusion
Acknowledgements
We sincerely thank the reviewers for their thoughtful feedback, insightful suggestions, and encouraging remarks, which helped improve the quality of this work. This research was supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC). Md Mahadi Hasan Nahid was supported by the Alberta Innovates Graduate Student Scholarship. This work was conducted during an internship at Huawei Technologies Canada Co., Ltd., Burnaby, Canada.
@@ Appendix B Error Analysis
These error types reflect both lexical and semantic challenges in schema linking, and they motivate future work on enhanced disambiguation, context-aware retrieval, and reasoning-aware schema grounding. Table 14 summarizes these error categories and their relative frequencies and some examples are discussed in Appendix H.
@@
Appendix C Implementation Details
@@
Appendix D Design Rationale and Justification
@@
Appendix E Extended Related Work
@@
Approaches like E-SQL caferoğlu2025esqldirectschemalinking enrich questions with schema context to enable direct alignment, while RSL-SQL Cao et al. (2024) utilizes partial query execution to validate links and improve robustness.
@@
The advent of LLMs shifted the paradigm toward in-context and retrieval-based promptingLee et al. (2025); Liu et al. (2025); Nahid and Rafiei (2025). MCS-SQL Lee et al. (2025) shows that using multiple prompts and multiple-choice ranking significantly improves accuracy by broadening the SQL generation search space. Solid-SQL Liu et al. (2024a) combines schema-aware prompting with retrieval strategies to improve robustness.
@@
Appendix F Schema Linking Examples
@@
Appendix G Prompt Templates
@@
Appendix H Representative Error Case